Sensical

The current Sensical service status is as follows :-

Network Web Hosting Email DNS Cloud
Green Green Green Green Green

There are no current open incidents. If you have an issue, please email our Support Team at support@sensical.net or, if a priority or you are without email access, please call +44 (0)844 686 6622. Please remember to quote your Sensical account name in any email.

The following incidents were closed in the last 30 days :-

Incident ID#60 opened on Sat 3 Feb 2024 11:26:20 +0000(GMT)
Current Service Impact:   Closed.

Update 3 Feb 24 16:38

DNS changes are now permitted as usual.
This incident is now closed. Relevent customers will receive a full incident report.
Some services will be failed back to primary datacentre in due course and this will be carried out under business-as-usual processes.
We apologise for any impact this incident caused you.

Status Change:   Incident closed.

Update 3 Feb 24 14:38

All network and web hosting services have been restored. DNS service is operating however DNS changes are not currently permitted. This will be resolved shortly after some further testing.

Revised Service Impact:   Net Green   Web Green   DNS Minor Issues  

Update 3 Feb 24 13:34

Power has been restored and non-geographically resilient services are coming back online.

Revised Service Impact:   Net Minor Issues   Web Minor Issues  

Update 3 Feb 24 12:33

Update - The power feeds affected at the LN6 datacentre are A1 and B1. C1 is still operational. Work is underway to restore power but we do not have an ETA at this time. All immediatel failover plans have been completed. Customers with geo-DR services are all now active.

Update 3 Feb 24 11:50

Power failure at LN6 datacentre confirmed. This outage coincides with a non service affecting scheduled maintenance taking place to upgrade the Engine Generator Paralleling System. This should not have impacted services. We are awaiting an official statement from datacentre although our team are in constant communication with the on-site team. Services with manual failver processes are in process of being failed over to the backup DC.

3 Feb 24 11:26

Issues at the LN6 datacentre are affecting web hosting. Investigating.

Service Impact:   Net Issues   Web Issues  

Admin